# Can a Multichoice Dataset be Repurposed for Extractive Question Answering?

### v1

@@ Abstract
The rapid evolution of Natural Language Processing (NLP) has favoured major languages such as English, leaving a significant gap for many others due to limited resources. This is especially evident in the context of data annotation, a task whose importance cannot be underestimated, but which is time-consuming and costly. Thus, any dataset for resource-poor languages is precious, in particular when it is task-specific. Here, we explore the feasibility of repurposing existing datasets for a new NLP task: we repurposed the Belebele dataset  Bandarkar et al. (2023), which was designed for multiple-choice question answering (MCQA), to enable extractive QA (EQA) in the style of machine reading comprehension. We present annotation guidelines and a parallel EQA dataset for English and Modern Standard Arabic (MSA). We also present QA evaluation results for several monolingual and cross-lingual QA pairs including English, MSA, and five Arabic dialects. We aim to enable others to adapt our approach for the 120+ other language variants in Belebele, many of which are deemed under-resourced. We also conduct a thorough analysis and share our insights from the process, which we hope will contribute to a deeper understanding of the challenges and the opportunities associated with task reformulation in NLP research.
Can a Multichoice Dataset be Repurposed
for Extractive Question Answering?
Teresa Lynn,1 Malik H. Altakrori,2 Samar M. Magdy1 Rocktim Jyoti Das,1 Chenyang Lyu,1 Mohamed Nasr,2 Younes Samih,2 Alham Fikri Aji,1 Preslav Nakov,1 Shantanu Godbole,2 Salim Roukos,2 Radu Florian,2 Nizar Habash3 1Mohamed bin Zayed University of Artificial Intelligence 2IBM Research AI  3New York University Abu Dhabi teresa.lynn@mbzuai.ac.ae, malik.altakrori@ibm.com, nizar.habash@nyu.edu
@@ 1 Introduction
Recent years have brought very fast developments in Natural Language Processing (NLP). However, this progress has not been equal for all languages, and most research has focused on English and a handful of other languages, with the vast majority of other languages being overlooked Joshi et al. (2020); Rehm and Way (2023). This is mainly due to the lack of data resources, which hampers the development of NLP tools for these languages. While resources have been developed for some languages, they are often for very specific tasks, and for very specific formulations of these tasks. With this in mind, here we explore the idea of re-purposing an existing dataset for low-resourced languages for a new task.
Starting from a motivation of creating QA datasets for Dialectal Arabic (DA)-MSA, we explore the possibility of repurposing the Belebele dataset Bandarkar et al. (2023), a multilingual dataset, which was created for the task of multiple-choice question answering, to be suitable for extractive question answering in the style of machine reading comprehension.
In the field of Question Answering, there are a number of different ways to answer a question  Wang (2022), e.g., (a) multiple-choice (MCQA) with answer choices, and context provided, (b) extractive (EQA) with context provided, (c) abstractive with context provided, and (d) open-ended with no context provided. An interesting research question is whether or not we can traverse between them, i.e., if we have a resource for one, can we repurpose it easily for a different reformulation of the task? Here we study this research question, looking into transferring from MCQA to EQA. Our contributions can be summarized as follows:
@@
  • We provide a dialectal Arabic QA benchmark by evaluating our EQA system’s performance on monolingual and cross-lingual QA pairs of {Eng, MSA}-{Eng, MSA, five DA}.
  • We create useful guidelines and a solid starting point for those interested in working on EQA for the other Belebele languages.
@@ 2 Related Work
We review related work on multilingual question-answering datasets and efforts to convert datasets between different formats, as well as present some relevant background on Arabic and its dialects.
@@ 2.1 Multilingual QA Datasets
Numerous multilingual QA datasets have already been developed e.g., XQUAD Artetxe et al. (2020), TyDi QA Clark et al. (2020), and MLQA Lewis et al. (2020), which focus on extractive-based QA. The recent ArabicaQA dataset Abdallah et al. (2024) represents a significant leap forward, being the first large-scale dataset specifically designed for Arabic QA. ArabicaQA’s extensive collection of over 89,000 questions, derived from a diverse set of Standard Arabic documents, provides a robust platform for advancing Arabic NLP, particularly in the context of large-language models. Additionally, alternative formats of multilingual QA have been explored, including open-ended QA with Mintaka Sen et al. (2022) and MCQA with Belebele Bandarkar et al. (2023). With the exception of Belebele, these datasets encompass a relatively narrow range of languages and do not include Arabic dialects. To address this gap, we propose leveraging the Belebele dataset, by including its dialectal content into a new parallel extractive-based QA format.
@@ 2.2 Converting MCQA Datasets
Some attempts have been made to convert an existing dataset into a different format. Specifically for QA, there have been conversion of MCQA datasets into natural language inference (NLI) Demszky et al. (2018); Khot et al. (2018). In these works, the premise is taken from the context paragraph, the correct choice is used as an entailment label, and the other choices are used as neutral or contradictory. Both approaches require effort to rephrase the question sentence into a statement, e.g., using rule-based methods Demszky et al. (2018) or human annotation Khot et al. (2018). Hadifar et al. (2022) introduce a new dataset specifically designed for the educational setting, focusing on converting between multiple-choice and open ended question formats. Conversely, automatically generating MCQA has also been explored Mitkov and Ha (2003); Kurdi et al. (2020); Ai et al. (2015); Karamanis et al. (2006), where the challenge involved generating distractor choices.
To our knowledge, there is no previous work on the automatic transformation of MCQA to EQA datasets. In this paper, we investigate the feasibility of such a task with an eye towards future automation of the process.
Table 1: English examples demonstrating some interesting types of MCQs in the Belebele dataset, and the respective issues they pose for extractive QA in this work. The Arabic version is shown in Appendix B.
@@ 2.3 A Note on Arabic and its Dialects
The Arabic language is a family of variants, among which Modern Standard Arabic (MSA) is the Arab world shared language of culture, media, and education. However, MSA is not the native language of any speaker of Arabic, whose day-to-day language is typically speak local variety, i.e., Dialectal Arabic (DA) that can be quite different from MSA and other dialects Salameh et al. (2018). MSA and DA coexist in what is called a diglossic relationship where different variants are used in different contexts Ferguson (1959). In the context of Arabic QA, there are many efforts that focus primarily on MSA Shaheen and Ezzeldin (2014); Mozannar et al. (2019); Biltawi et al. (2021); Alwaneen et al. (2022). A recent exception is the work of Faisal et al. (2021) which works on spoken dialectal QA and includes Arabic among other languages.
The common wisdom in Arabic NLP is that systems should be robust to DA forms, but generate output primarily in MSA. This is reasonable for QA in particular, given the relatively larger trusted content in MSA compared to that which dominates social media. As such, in this paper, we focus on modeling QA that accepts questions in a range of dialects and answers in standard languages (MSA and EN). We select five representative dialects from across the Arab world to work with: Egyptian (arz), Iraqi (acm), Moroccan (ary), Gulf (Najdi; ars), and Levantine (North; apc).
@@ 3.1 Belebele Dataset
Belebele Bandarkar et al. (2023) is an open-source multiple-choice machine reading comprehension (MRC) dataset that covers 122 languages and language variants, including English, MSA, and a number of Arabic dialects. Each question is based on a short passage from the Flores-200 dataset Team et al. (2022) and has four multiple-choice answers, with the correct answer identified. This benchmark dataset enabled the evaluation of natural language understanding (NLU) across high/medium/low-resource languages and language variants.
With respect to the style of the dataset content, the creators of the Belebele dataset indicated that they selected multiple-choice questions (MCQs) because it would lead to the fairest evaluation across languages and MCQs enable the ability to scale to many languages when translating from English, an important design goal for Belebele. They worked with a Language Service Provider (LSP) to create the questions and the multi-choice answers in an iterative process, by developing guidelines that included rules such as no use of double negatives, only use answers that are decently plausible to require the test-taker to fully read and understand the passage Bandarkar et al. (2023).
Since our effort is part of a larger project on Arabic QA, we deemed the Belebele dataset to be a suitable starting point for establishing an EQA benchmark for Arabic dialect QA when paired with MSA and English passages. Our five dialects of focus represent the full spectrum of variations across the Arab world: Gulf, Levantine, Iraqi, Egyptian, and Moroccan. We have questions in English, MSA and Arabic dialects, with the answers in MSA and English only.
@@ 3.2 The MCQ Zoo
When examining the different Belebele MCQs and their associated passages, we recognized a considerable number of MCQ types that vary on a continuum from simple word matching to required passage and world understanding. Table 1 presents a set of English examples along this continuum to highlight the challenges in identifying text spans in the associated passages. The Arabic version is in Appendix B.
The simplest types are perhaps Exact Match, Partial Overlap Match and Morphological Match, Table 1.(a, b, c), respectively, where the MCQ answer is “present” in the passage. Relatively more complex are cases that may be resolvable through Paraphrasing, Entailment, or Semantic Similarity, in Table 1.(d, e, f), respectively. Next comes the cases that require higher levels of awareness of context, from that of Coreference in the text all the way to Pragmatic Knowledge, in Table 1.(g-i). Finally we present examples of complex kinds of MCQ, where wording of the question is highly dependent on the provided multi-choice answers. While a span can be identified in principle in Table 1.(j), the question may not stand alone with the passage in Table 1.(k, l).
These challenges made it apparent that (a) not all MCQAs are usable for the EQA task; and (b) that human annotation was a more realistic approach than automating the conversion of the dataset. While the first few MCQ types suggest that an automatic process could be used to identify the QA span, the rest, in increasingly difficulty, suggest that the MCQ answers are not as helpful. The last two types are not usable in our assessment for this task. We took all of these insights into consideration as we developed the guidelines and annotated the spans, which we discuss next.
@@ 4.3 Annotation Guidelines
Informed by the findings of the initial pilot study, we created two separate versions of the guidelines for English and Arabic, accounting for linguistic differences that influenced the wording of the instructions and the need for language-specific examples. See Appendices C and D for the complete EN and MSA guidelines, respectively.
@@
Annotators were also given clear guidance on what to do with articles, prepositions, punctuation, and instances where the answer appeared more than once in the passage. Equipped with better awareness of the issues described in Section 3.2, for each QA pair, the annotators could either (i) annotate the span and if necessary, label the QA as ‘BB’ to indicate it as being complex and problematic, as per Table 1.(d,e,g,i) or (ii) label the QA as ‘X’ to be removed due to the answer was impossible to find, as per Table 1.(k,l).
Finally, annotators were advised that it was more important to find a relevant span (in which evidence of the answer could be found) than finding an exact span that may not exist. This guidance was a significant help given the many differences between the nature of MCQA and EQA.
@@ 4.4 English Data Annotation
We took a bootstrapping approach to annotation across the two languages. Given the parallel nature of the dataset, the EN annotation was carried out first, followed by a semi-automatic approach to annotating the MSA data (Section 4.5). The annotations were done on the remaining 415 QA pairs in the Belebele dataset (including a review and correction of the first 100 from the pilot round). Eleven annotators, who were either fluent or native English speaking, annotated the QA pairs, with the annotations subsequently reviewed by independent reviewers.
The distribution of the QA pairs was as follows: 50 QA pairs were used in our IAA study (see Section 4.6) and were annotated by three annotators. The remaining 365 pairs were distributed amongst the other eight annotators, where six annotators had 50 pairs each, and the remaining two annotators had 42 and 23, respectively. The annotators were presented with the English Belebele passage, the original English Belebele question and a copy of the passage in which they would mark the span of text in which the answer could be found. They also had the option of including comments for problematic instances that required broader discussion or verification.
As a result of the annotation process, 86 out of the 415 QA pairs were labelled as ‘X’ and excluded. Out of the remaining 329 pairs, 45 were labelled as ‘BB’. These BB QAs remain part of the final dataset, as we are also interested in establishing what difficulties they pose for a QA system, given the need for reasoning and resolving linguistic phenomena like paraphrasing, pragmatics and coreference resolution.
@@ 4.5 Arabic Data Annotation
    Marking the same questions in MSA that were labelled as ’X’ in English so that they are not reviewed by annotators.
@@
    Identifying the MSA span that has the same length and exhibits the highest n-gram overlap with the EN span. A sliding-window approach was used.
@@
    If no equivalent translation was found, an approximation of location in the MSA passage was used based on the EN span’s word off-set. [^2] An example of this case is when the span is 1-2 words long and its translation is not found in the MSA passage.
@@
    Native Arabic speakers reviewed the pre-annotated spans according to the MSA-based annotation guidelines.
The 415 QA pairs were divided amongst four native Arabic speaking annotators, with the annotations subsequently reviewed by independent reviewers. The annotators were presented with the MSA Belebele passage, the original MSA Belebele question and a copy of the passage where the span was pre-annotated. During the review process, the annotators could either modify the span markers, and label as ‘BB’ where necessary, or label the QA as ‘X’ to be removed.

### v2

# From Multiple-Choice to Extractive QA: A Case Study for English and Arabic

@@ Abstract
The rapid evolution of Natural Language Processing (NLP) has favoured major languages such as English, leaving a significant gap for many others due to limited resources. This is especially evident in the context of data annotation, a task whose importance cannot be underestimated, but which is time-consuming and costly. Thus, any dataset for resource-poor languages is precious, in particular when it is task-specific. Here, we explore the feasibility of repurposing an existing multilingual dataset for a new NLP task: we repurpose a subset of the Belebele dataset  Bandarkar et al. (2023), which was designed for multiple-choice question answering (MCQA), to enable the more practical task of extractive QA (EQA) in the style of machine reading comprehension. We present annotation guidelines and a parallel EQA dataset for English and Modern Standard Arabic (MSA). We also present QA evaluation results for several monolingual and cross-lingual QA pairs including English, MSA, and five Arabic dialects. We aim to help others adapt our approach for the remaining 120 Belebele language variants, many of which are deemed under-resourced. We also provide a thorough analysis and share insights to deepen understanding of the challenges and opportunities in NLP task reformulation.
From Multiple-Choice to Extractive QA:
A Case Study for English and Arabic
Teresa Lynn,1 Malik H. Altakrori,2 Samar M. Magdy1 Rocktim Jyoti Das,1 Chenyang Lyu,1 Mohamed Nasr,2 Younes Samih,2 Kirill Chirkunov,1 Alham Fikri Aji,1 Preslav Nakov,1 Shantanu Godbole,2 Salim Roukos,2 Radu Florian,2 Nizar Habash1,3 1Mohamed bin Zayed University of Artificial Intelligence 2IBM Research AI  3New York University Abu Dhabi teresa.lynn@mbzuai.ac.ae, malik.altakrori@ibm.com, nizar.habash@nyu.edu
@@ 1 Introduction
Recent years have brought about very fast developments in Natural Language Processing (NLP). However, this progress has not been equal for all languages, and most research has focused on English and a handful of other languages, with the vast majority of other languages being overlooked Joshi et al. (2020); Rehm and Way (2023). This is mainly due to the lack of data resources, which hampers the development of NLP tools for these languages. While many resources have been developed for some languages, they are often for very specific tasks, and for very specific formulations of these tasks.
Starting from a motivation of creating cross-lingual Question Answering (QA) datasets for an under-resourced pair, Dialectal Arabic (DA) – Modern Standard Arabic (MSA), we explore the possibility of re-purposing the multilingual Belebele dataset Bandarkar et al. (2023), which was created for the task of multiple-choice question answering, to be suitable for extractive question answering in the style of machine reading comprehension.
In the field of Question Answering, there are a number of different ways to answer a question  Wang (2022), e.g., (a) multiple-choice QA (MCQA) with answer choices, and context provided, (b) extractive QA (EQA) with context provided, (c) abstractive QA with context provided, and (d) open-ended QA with no context provided. An interesting research question is whether we can traverse between them, i.e., if we have a resource for one QA task, can we repurpose it easily for a different reformulation of the task? Here we study this research question, in the context of transferring from MCQA to EQA, as we deem EQA more useful in real-world applications where the correct answer is often not known. Our contributions are:
@@
  • We provide a dialectal Arabic QA benchmark by evaluating our EQA system’s performance on monolingual and cross-lingual QA pairs of {EN, MSA}–{EN, MSA, five DAs}.
  • We provide guidelines and a strong foundation for creating EQA datasets for the remaining 120 Belebele languages.
@@ 2.1 Multilingual QA Datasets
Numerous multilingual QA datasets have already been developed e.g., XQUAD Artetxe et al. (2020), TyDi QA Clark et al. (2020), and MLQA Lewis et al. (2020), which focus on extractive-based QA. The recent ArabicaQA dataset Abdallah et al. (2024) represents a significant leap forward, being the first large-scale dataset specifically designed for Arabic QA: it has over 89k questions, derived from a diverse set of Standard Arabic documents, providing a robust platform for advancing Arabic NLP, particularly in the context of large language models. Additionally, alternative formats of multilingual QA have been explored, including open-ended QA with Mintaka Sen et al. (2022) and MCQA with Belebele Bandarkar et al. (2023). With the exception of Belebele, these datasets encompass a relatively narrow range of languages and do not include Arabic dialects. To address this gap, we propose leveraging the Belebele dataset, to enable the inclusion of its dialectal content into a new parallel DA – MSA/EN dataset.
@@ 2.2 Converting MCQA Datasets
Some attempts have been made to convert existing datasets into a different format. Specifically for QA, there has been conversion of MCQA datasets into natural language inference (NLI) Demszky et al. (2018); Khot et al. (2018): the premise is taken from the context paragraph, the correct choice is used as an entailment label, and the other choices are used as neutral or contradictory. Both approaches require effort to rephrase the question into a statement, e.g., using rule-based methods Demszky et al. (2018) or human annotation Khot et al. (2018). Hadifar et al. (2023) introduced a new dataset specifically designed for the educational setting, focusing on converting between multiple-choice and open-ended question formats. Automatically generating MCQA has also been explored Mitkov and Ha (2003); Kurdi et al. (2020); Ai et al. (2015); Karamanis et al. (2006), where the challenge is to generate distractor choices.
To our knowledge, there is no previous work on converting MCQA to EQA datasets. In this paper, we investigate the feasibility of such a task with an eye towards future automation of the process.
Table 1: English examples demonstrating some interesting types of MCQs in the Belebele dataset, and the respective issues they pose for extractive QA in this work. The Arabic version is shown in Appendix C.
@@ 2.3 A Note on Arabic and its Dialects
The Arabic language is a family of variants, among which Modern Standard Arabic (MSA) is the Arab world’s shared language of culture, media, and education. However, MSA is not the native language of any speaker of Arabic, whose day-to-day language is typically a local variety, i.e., Dialectal Arabic (DA) that can be quite different from MSA and other dialects Salameh et al. (2018). MSA and DA coexist in what is called a diglossic relationship where different variants are used in different contexts Ferguson (1959). In the context of Arabic QA, there are many efforts that focus primarily on MSA Shaheen and Ezzeldin (2014); Mozannar et al. (2019); Biltawi et al. (2021); Alwaneen et al. (2022). A recent exception is the work of Faisal et al. (2021), which works on spoken dialectal QA and includes Arabic among other languages.
The common wisdom in Arabic NLP is that systems should be robust to handling DA forms, but should generate output primarily in MSA. This is reasonable for QA in particular, given the relatively larger trusted content in MSA compared to DA which dominates social media. As such, in this paper, we focus on modeling QA that accepts questions in a range of dialects and provides answers in standard languages (MSA and EN). We selected the following five representative dialects from across the Arab world to work with: Egyptian (arz), Iraqi (acm), Moroccan (ary), Gulf (Najdi; ars), and Levantine (North; apc).
@@ 3.1 Belebele Dataset
Belebele Bandarkar et al. (2023) is an open-source multiple-choice machine reading comprehension (MRC) dataset that covers 122 languages and language variants, including English, MSA, and five Arabic dialects. Each question is based on a short passage from the Flores-200 dataset Team et al. (2022) and has four multiple-choice answers, with the correct answer identified. This benchmark dataset enabled the evaluation of natural language understanding (NLU) across high/medium/low-resource languages and language variants.
With respect to the style of the dataset content, the creators of the Belebele dataset indicated that they selected multiple-choice questions (MCQs) because it would lead to the fairest evaluation across languages and MCQs enable the ability to scale to many languages when translating from English. They worked with a Language Service Provider to create the questions and the multi-choice answers in an iterative process, by developing guidelines that included rules such as no use of double negatives, only use answers that are decently plausible to require the test-taker to fully read and understand the passage Bandarkar et al. (2023).
Our effort is part of a larger project on Arabic QA; as such we deemed the Belebele dataset to be a suitable starting point for establishing an EQA benchmark for Arabic dialect QA when paired with MSA and English passages. Our five dialects of focus represent the full spectrum of variations across the Arab world: Gulf, Levantine, Iraqi, Egyptian, and Moroccan. We have questions in English, MSA and Arabic dialects, with the answers in MSA and English only.
@@ 3.2 The MCQ Zoo
When examining the different Belebele MCQs and their associated passages, we recognized a considerable number of MCQ types that vary on a continuum from simple word matching to required passage and world understanding. Table 1 presents a set of English examples along this continuum to highlight the challenges that hindered us from employing a fully automated process in identifying text spans in the associated passages. The Arabic version is in Appendix C.
The simplest MCQ types are perhaps Exact Match, Partial Overlap Match and Morphological Match, Table 1.(a, b, c), respectively, where the MCQ answer is “present” in the passage. Relatively more complex are cases that may be resolvable through Paraphrasing, Entailment, or Semantic Similarity, in Table 1.(d, e, f), respectively. Next comes the cases that require higher levels of awareness of context, from that of Coreference in the text all the way to Pragmatic Knowledge, in Table 1.(g-i). Finally we present examples of complex kinds of MCQ, where wording of the question is highly dependent on the provided multi-choice answers. While a span can be identified in principle in Table 1.(j), the question may not stand alone with the passage in Table 1.(k, l).
These challenges made it apparent that (a) not all MCQAs are usable for the EQA task; and (b) that human annotation was a more realistic approach than automating the conversion of the dataset. While the first few MCQ types suggest that an automatic process could be used to identify the QA span, the rest, in increasingly difficulty, show that the MCQ answers are not as helpful. The last two types are not usable in our assessment for this task. We took all of these insights into consideration as we developed the guidelines and annotated the spans, which we discuss next.
@@ 4.3 Annotation Guidelines
Informed by the findings of the initial pilot study, we created two separate versions of the guidelines for English and Arabic, accounting for linguistic differences that influenced the wording of the instructions and the need for language-specific examples. See Appendices D and E for the complete EN and MSA guidelines, respectively.
@@
Annotators were also given clear guidance on what to do with articles, prepositions, punctuation, and instances where the answer appeared more than once in the passage (Appendices D and E). Equipped with better awareness of the issues described in Section 3.2, for each QA pair, the annotators could either: (i) annotate the span and if necessary, label the QA as ‘BB’ to indicate it as being complex and problematic, as per Table 1.(d,e,g,i) or (ii) label the QA as ‘X’ to be removed, due to the answer being impossible to find, as per Table 1.(k,l). Finally, annotators were advised that it was more important to find a relevant span (in which evidence of the answer could be found) than finding an exact span that may not exist. This guidance was a significant help given the many differences between the nature of MCQA and EQA.
@@ 4.4 English Data Annotation
We took a bootstrapping approach to annotation across the two languages. Given the parallel nature of the dataset, the EN annotation was carried out first, followed by a semi-automatic approach to annotating the MSA data (Section 4.5). The annotations were done on the remaining 415 QA pairs in the Belebele dataset (including a review and correction of the first 100 from the pilot round). Eleven annotators, both fluent and native English speakers, annotated the QA pairs, with the annotations subsequently reviewed by independent reviewers.
The distribution of the QA pairs was as follows: 50 QA pairs were used in our IAA study (see Section 4.6) and were annotated by three annotators. The remaining 365 pairs were distributed amongst the other eight annotators, where six annotators had 50 pairs each, and the remaining two annotators had 42 and 23, respectively. The annotators were given the English Belebele passage, the original question and a copy of the passage to mark the answer span. They could also add comments for problematic instances that required broader discussion or verification.
As a result of the annotation process, 86 out of the 415 QA pairs were labelled as ‘X’ and excluded. Out of the remaining 329 pairs, 44 were labelled as ‘BB’. These BB QAs remain part of the final dataset, as we are also interested in establishing what difficulties they pose for a QA system, given the need for reasoning and resolving linguistic phenomena like paraphrasing, pragmatics and coreference resolution.
@@ 4.5 Arabic Data Annotation
    Marking as ‘X’ the MSA questions labelled ‘X’ in English to avoid re-annotation. In our experience, this approach proved consistent for both English and MSA.
@@
    Using a sliding-window approach to identify the MSA span with the same length and highest n-gram overlap with the EN span.
@@
    If no equivalent translation was found, an approximate MSA location based on the EN span’s word offset was used. [^2] An example of this case is when the span is 1-2 words long and its translation is not found in the MSA passage.
@@
    Native Arabic speakers reviewed the pre-annotated spans according to the MSA-based annotation guidelines and made corrections where necessary.
The 415 QA pairs were divided amongst four native Arabic speaking annotators, with the annotations subsequently reviewed by independent reviewers. The annotators were presented with the MSA Belebele passage, the original MSA Belebele question and a copy of the passage where the span was pre-annotated. During the review process, the annotators could either: leave the span unchanged or modify the span markers, and label as ‘BB’ where necessary, or label the QA as ‘X’ to be removed.